# Neural Signals Generate Clinical Notes in the Wild

### 06 Mar 2026 v2

@@ 4.3 Sequence-Aware Alignment
To enable EEG-conditioned clinical report generation and leverage LLMs’ generative capabilities, EEG tokens must be aligned with the text embedding space of LLMs. In the vision-language setting, models such as LLaVA achieve this via a simple linear projection (Liu et al., 2023). However, this paradigm does not generalize well to EEG (Section 5.4) due to its inherent sequential nature, unlike static images. EEG signals exhibit long-range temporal dynamics across epochs that must be preserved during alignment. As a result, we face two key challenges: modeling long temporal dependencies and scalability for very long sequences. These challenges motivate the design of two alignment strategies in the proposed Sequence-Aware Alignment module (see Figure 1b.2): (1) Sequence Context Alignment, which preserves the full temporal sequence, and (2) Sequence Context Compression, which compresses the sequence into a fixed-length latent representation for memory-efficient alignment.
@@
Given a sequence of epoch tokens $\mathbf{E}_{\text{eeg}}$, we augment the sequence with learnable position encodings and [SESSION_SEP] tokens to differentiate EEG sessions. A lightweight linear-attention transformer (Wang et al., 2020; Katharopoulos et al., 2020) is then applied to capture the temporal structure across epochs. The resulting representations are subsequently projected through a linear layer into the language embedding space, yielding $\mathbf{H}_{\text{eeg}}\in\mathbb{R}^{N\times D_{\text{llm}}}$, where $D_{\text{llm}}$ denotes the LLM embedding dimension.
@@
In an attempt to tackle the memory constraints and improve scalability for long EEG sequences, we explore a compression strategy inspired by Perceiver architectures (Jaegle et al., 2021b, a). After encoding temporal structure among epoch tokens, a fixed set of learnable query tokens $\mathbf{Q}\in\mathbb{R}^{L\times D_{\text{eeg}}}(L<N)$ attend to the full sequence of epoch tokens via cross-attention. This compresses the variable-length EEG sequence into a fixed number of tokens, which are subsequently projected into the language embedding space, enabling efficient alignment for recordings of arbitrary duration.
@@ 4.4 Prompt Fusion and Generation
CELM generates clinical reports by conditioning a local LLM on EEG-derived representations jointly with clinical context when available (see Figure 1b.3). Given projected EEG tokens $\mathbf{H}_{\text{eeg}}$, we augment the sequence with special tokens [EEG_START] and [EEG_END] to distinguish EEG representations from text inputs. A task-specific prompt that specifies the target report sections, together with optional clinical context (e.g., reason for study or clinical history), is tokenized using the LLM tokenizer to obtain text embeddings $\mathbf{H}_{\text{prompt}}$. The final input sequence to the LLM is formed by concatenation,
@@ 5 Experiments and Results
Table 1: Report generation performance comparison on samples with clinical context across two hospital sites (S0001 and S0002). We evaluate BLEU-1, ROUGE-1, and METEOR under two input settings: (i) Unimodal + Text Only Input, where language models generate reports solely from clinical context text, and (ii) Unimodal + Text + EEG Features Input, where EEG-derived features are additionally provided. The table compares multiple strong baselines, including general-purpose and medical LLMs, against our CELM . The best results are highlighted in orange, and the best baseline per category is highlighted in blue. Relative improvements are also reported.
[…]
@@
S0002: From the available visit types (EEG, long-term monitoring, portable EEG), we retained only standard EEG visits with single corresponding sessions, yielding $4,873$ reports from $4,379$ patients.
@@
As no existing EEG multimodal models with generative capabilities are available for EEG report generation, we establish baselines by augmenting local LLMs with EEG-derived inputs. We evaluate two settings: (1) Unimodal + text-only input, where the LLMs receive only clinical context as input without EEG data, providing a lower bound that quantifies hallucination and measures performance when the model does not attend to neural signals, and (2) Unimodal + text and EEG features input, where we augment the clinical context with channel-wise spectral band power features (delta, theta, alpha, beta, gamma) extracted from each EEG session, addressing context length limitation that prevents providing raw EEG arrays directly. We evaluate LLMs from three model families, including LLaMA-3 (Grattafiori et al., 2024), Gemma-3 (Team et al., 2025), and Qwen-3 (Yang et al., 2025), across multiple parameter scales. Additionally, we include MedGemma (Sellergren et al., 2025) as a medical LLM baseline. Additional implementation details are provided in Appendix B. We report smoothed BLEU-1 (Papineni et al., 2002), ROUGE-1 (Lin, 2004), and METEOR (Banerjee and Lavie, 2005) scores to evaluate report generation performance. Comprehensive results and metrics are provided in Appendix D.
5.1 Report Generation Performance Comparison
Table 1 compares CELM against unimodal baselines using either text-only or text with handcrafted EEG features as input, evaluated on test samples containing patient history. We report results for two CELM variants: CELM-SCC, which uses sequence context compression to compress and produce a fixed number of projected EEG tokens, and CELM , which uses non-compressed EEG representations via sequence context alignment. When EEG features are provided, most baseline LLMs show modest improvements over text-only inputs on S0002 ($0.1912\rightarrow 0.2483$), except LLaMA-3.1-8B, which fails due to its 8K token context limit. However, on S0001, performance degrades across baselines due to greater report complexity introduced by multiple sections.
Both CELM variants substantially outperform all baselines across all metrics, with CELM achieving the best results. On average, CELM yields a two-fold improvement over the strongest baseline across both sites and all evaluation metrics. For example, on the ROUGE score for S0002, performance improves from $0.2886$ to $0.6408$ relative to the strongest baseline. Although CELM-SCC outperforms the baselines, its performance gap relative to the non-compressed model ($0.4487$ vs $0.6408$) underscores an important open challenge. This leads to an important future direction of developing efficient compression strategies that preserve clinically relevant information from long EEG sequences while remaining within LLM context constraints.
Table 2: Zero-context report generation performance on S0002. The best results are highlighted in orange, and the best baseline is highlighted in blue. Relative improvements are reported with respect to the strongest baseline under each setting.
@@ 5.1 Report Generation Performance Comparison
Figure 3: (a) Report generation performance of different projector variants. (b) Training dynamics of each variant, including training loss, validation loss, and perplexity curves.
@@ 5.2 Zero-Context Report Generation
To assess our model’s ability to generate clinical reports solely from EEG signals, without relying on clinical context, we define a zero-context report generation task in which only EEG data is provided as input. This experiment also serves as a test of potential language-model bias, in which reports could be generated by inferring plausible clinical context while ignoring EEG inputs. We evaluate this task on a subset of the S0002 dataset containing reports without patient clinical context, and the results are summarized in Table 2. CELM significantly outperforms all baselines across all metrics, demonstrating that it effectively extracts and utilizes clinically relevant information from EEG recordings.
5.3 Performance Analysis by Report Section
Clinical EEG reports comprise multiple sections, each serving a distinct purpose: EEG description/details provides a detailed narrative of observed waveforms and patterns, impression/interpretation summarizes the clinical significance of findings, background activity characterizes baseline rhythms, and events/seizures documents seizure episodes. To evaluate whether CELM can reliably generate these diverse sections, we conduct a section-wise analysis on S0001, which exhibits richer section diversity compared to S0002 (where reports predominantly contain only EEG description/details; see Figure 5 in Appendix). Figure 2 presents ROUGE-1 scores, comparing CELM against the best-performing unimodal baseline (text + EEG features) from each LLM family. CELM achieves the highest performance in 6 out of 7 report sections. Performance degrades in the interictal epileptiform abnormalities section, highlighting a key limitation and a challenge for ELMs in modeling rare and clinically complex events, an important direction for future work. Detailed results are provided in Table 12.
@@ 5.4 Alignment Module Ablation
The alignment module is a critical component of EEG-language models. In our initial experiments, we found that naive linear projection into the LLM embedding space is insufficient for the EEG domain. In this section, we present an ablation study comparing different alignment strategies. We evaluate two baseline alignment modules: (1) Linear projector, which directly maps EEG embeddings to the LLM embedding space using a single linear layer, and (2) Perceiver projector, which employs learnable query tokens that attend to EEG sequences via cross-attention followed by feedforward and projection. We compare these against our proposed SCA and SCC projectors, described in Section 4.3. Results are summarized in Figure 3.
As shown in Figure 3a, the SCA projector consistently achieves the best performance across all evaluation metrics. Analysis of training dynamics in Figure 3b further highlights the importance of modeling inter-epoch temporal dependencies prior to alignment. While the Linear and Perceiver projectors’ training losses decrease, their validation losses and perplexities increase, indicating overfitting. In contrast, both SCA and SCC exhibit more stable convergence, with lower validation loss and perplexity over time, and SCA converges faster across all variants. Notably, SCA achieves lower validation perplexity than other variants, indicating substantially higher confidence in token generation. Overall, this study confirms that modeling temporal dependencies between EEG epochs before projection is essential for effective EEG-to-language alignment.
5.5 Qualitative Analysis
Figure 4 presents qualitative comparisons on S0002 between the best-performing unimodal baseline (text with EEG features), the linear projector variant, and CELM , alongside ground truth reports. Figure 4a shows the EEG description/details section, while Figure 4b shows the impression/interpretation section. Across these examples, CELM produces reports that are more closely aligned with the ground truth, correctly identifying clinically relevant findings such as moderate bilateral beta activity, intermittent drowsiness, and bilateral slowing in the theta and delta ranges. In contrast, the linear projector and unimodal baselines fail to capture these. In the impression section, CELM identifies the recording as abnormal, whereas the linear projector variant incorrectly predicts it as normal. These results demonstrate the promise of ELMs in learning to generate clinical reports from unstructured notes paired with long EEG recordings. At the same time, they underscore the need for more rigorous and standardized benchmarking to assess these models. Additional examples spanning both sites, diverse report sections, and varying performance levels are provided in Appendix D.7 to further illustrate model strengths and failure modes.
6 Discussion and Implication
@@ B.3 Baselines and Prompts
In this section, we provide additional details on the baseline models and the prompts used in our experiments. Table 4 summarizes all LLMs considered in our study along with their respective context lengths [^1] Context length information is extracted from their respective technical reports and Hugging Face https://huggingface.co/, further underscoring the practical challenges of directly conditioning LLMs on long-duration EEG recordings. Figures 6 and 7 present the prompts used for the unimodal text-only and unimodal text + EEG feature baselines, respectively. For clarity, we provide examples for certain prompt sections using bounding boxes.
Table 4: Summary of LLM Baselines
LLM Family Context Length
Gemma-3-1b-it Gemma 3 32K
Gemma-3n-E2B-it Gemma 3 32K
Gemma-3n-E4B-it Gemma 3 32K
Gemma-3-4b-it Gemma 3 128K
Medgemma-4b-it Gemma 128K
Llama-3.2-1B-Instruct Llama-3.2 128K
Llama-3.2-3B-Instruct Llama-3.2 128K
Llama-3.1-8B-Instruct Llama-3.1 128K
Meta-Llama-3-8B-Instruct Llama-3 8K
Qwen3-1.7B Qwen 3 32K
Qwen3-4B-Instruct-2507 Qwen 3 256K
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]

### 13 May 2026 v3

@@ 4.3 Sequence-Aware Alignment
To enable EEG-conditioned clinical report generation, EEG representations must be aligned with the text embedding space of LLMs. Existing vision-language models such as LLaVA achieve this through simple linear projection Liu et al. (2023), but this strategy does not generalize well to EEG (Section 5.5) due to its sequential and long-range temporal structure. Unlike static images, EEG signals contain temporal dynamics across epochs that must be preserved during alignment. This introduces two key challenges: modeling long temporal dependencies and scaling to very long sequences. To address these challenges, we propose two sequence-aware alignment strategies (Figure 3b): (1) Sequence Context Alignment, which preserves full temporal structure, and (2) Sequence Context Compression, which compresses EEGs into fixed-length latent representations for memory-efficient alignment.
@@
Given a sequence of tokens $\mathbf{E}_{\text{eeg}}$, we augment the sequence with position encodings and [SESSION_SEP] tokens to differentiate EEG sessions. A lightweight linear-attention transformer (Wang et al., 2020) is then applied to capture the temporal structure across epochs. The resulting representations are subsequently projected through a linear layer into the language embedding space, yielding $\mathbf{H}_{\text{eeg}}\in\mathbb{R}^{N\times D_{\text{llm}}}$, where $D_{\text{llm}}$ denotes the LLM embedding dimension.
@@
To address the memory and scalability challenges of long EEG sequences, we explore a compression strategy inspired by Perceiver architectures (Jaegle et al., 2021b, a). After encoding temporal structure among epoch tokens, a fixed set of learnable query tokens $\mathbf{Q}\in\mathbb{R}^{L\times D_{\text{eeg}}}(L<N)$ attend to the full sequence of epoch tokens via cross-attention. This compresses the variable-length EEG sequence into a fixed number of tokens, which are subsequently projected into the language embedding space, enabling efficient alignment for recordings of arbitrary duration.
@@ 4.4 Prompt Fusion and Generation
CELM generates clinical reports by conditioning a local LLM on EEG-derived representations jointly with clinical context when available (see Figure 3c). Given projected EEG tokens $\mathbf{H}_{\text{eeg}}$, we augment the sequence with special tokens [EEG_START] and [EEG_END] to distinguish EEG representations from text inputs. A task-specific prompt that specifies the target report sections, together with optional clinical context (e.g., reason for study or clinical history), is tokenized using the LLM tokenizer to obtain text embeddings $\mathbf{H}_{\text{prompt}}$. The final input sequence is formed by concatenation,
@@ 5 Experiments and Results
S0002: From the available visit types (EEG and portable EEG), we retained only standard EEG visits with single corresponding sessions, yielding $4,873$ reports from $4,379$ patients.
@@
Table 1: Report generation performance comparison on samples with clinical context across two hospital sites (S0001 and S0002). We report mean $\pm$ standard deviation for smoothed BLEU-1, ROUGE-1, and METEOR scores. The best results are highlighted in orange, and the best baseline per category is highlighted in blue. Relative improvements are also reported, and indicates finetuning.
[…]
@@
Due to the lack of existing EEG multimodal models for clinical EEG report generation, we compare CELM against three types of baselines: (1) Unimodal + text-only input: only clinical context is provided as input without EEG data, providing a lower bound that quantifies hallucination and isolates performance when the model does not attend to neural signals, (2) Unimodal + text and EEG features input: the clinical context is augmented with channel-wise spectral band power features (delta, theta, alpha, beta, gamma, Appendix B.1) extracted from each EEG session, overcoming context length limitation that prevents providing raw EEG arrays directly, (3) Multimodal + linear projector: a linear projector maps pretrained EEG embeddings from CBraMod Wang et al. (2024b) into the language model’s embedding space, enabling direct conditioning on continuous EEG signals. We evaluate LLMs from three model families, including LLaMA-3 (Grattafiori et al., 2024), Gemma-3 (Team et al., 2025), and Qwen-3 (Yang et al., 2025), across multiple parameter scales, and include MedGemma (Sellergren et al., 2025) as a domain-specific baseline (more details at Appendix B). We report lexical metrics Papineni et al. (2002); Lin (2004); Banerjee and Lavie (2005) along with manual qualitative analysis and a human expert evaluation study to evaluate report generation performance.
5.1 Report Generation Performance
Table 1 compares the report generation performance of CELM on test samples from both sites containing patient history. It compares against unimodal baselines that use either text-only or text with handcrafted EEG features as input, and finetuned multimodal baselines that use EEG foundation model embeddings. We report results for two CELM variants: CELM-SCC, which uses sequence context
Table 2: Zero-context report generation performance on S0001 and S0002. The best results are highlighted in orange, and the best baseline is highlighted in blue.
@@ 5.1 Report Generation Performance
compression to compress and produce a fixed number of projected EEG tokens, and CELM, which uses non-compressed EEG representations via sequence context alignment. Incorporating EEG features yields modest gains over text-only on S0002 ($0.1912\rightarrow 0.2483$) except for LLaMA-3.1-8B, which fails due to its 8K context limit. On S0001, performance degrades due to greater multi-section report complexity. Finetuned multimodal baselines, which align EEG foundation embeddings with the language space via linear projection, achieve the best baseline performance. However, the marginal gains relative to the additional training overhead underscore the fundamental challenge of EEG-language alignment.
Both CELM variants substantially outperform all baselines across metrics, with CELM achieving approximately two-fold improvement over the strongest baseline on both sites (e.g., ROUGE on S0002: $0.3084\rightarrow 0.6408$). Although CELM-SCC outperforms the baselines, its performance gap relative to the non-compressed model ($0.4487$ vs $0.6408$) highlights a key trade-off between compression and performance for ELMs. This leads to an important future direction: developing compression and alignment strategies that preserve clinically relevant information while scaling to longer EEG sequences.
Figure 4: Manual qualitative analysis across four clinically relevant dimensions: abnormality prediction, lateralization, spectral dynamics, and epileptiform abnormalities. Representative example cases are provided below each dimension.
@@ 5.2 Zero-Context Report Generation
To assess our model’s ability to generate clinical reports solely from EEG signals, without relying on clinical context, we define a zero-context report generation task in which only EEG data is provided as input. This experiment also studies the potential language-model bias, in which reports could be generated by inferring plausible clinical context while ignoring EEG inputs. We conduct this experiment on both sites, using only the EEG signal as input, and summarize the results in Table 2. CELM outperforms all baselines across all metrics, demonstrating that it effectively extracts and utilizes clinically relevant information from EEG recordings.
5.3 Manual Qualitative Analysis
While lexical metrics capture overlap, they fail to assess clinical correctness, and no reliable automated alternative exists. To address this, we manually inspect 30 randomly sampled reports from both sites and evaluate correctness across four clinically relevant dimensions. Results and representative examples are presented in Figure 4. ❐ Abnormality prediction. We evaluate whether the generated report correctly classifies the recording as normal/abnormal. CELM achieves higher accuracy than all baselines. However, it frequently misclassifies mildly abnormal recordings as normal. One important observation is that even when the model accurately describes abnormal signal patterns (e.g., “frequent bilateral but right-predominant temporal sharply contoured theta and delta slowing”), it often fails at decision-making. This suggests a lack of domain knowledge, in which injecting neurology guidelines presents a promising approach. ❐ Lateralization. We assess whether the generated report correctly identifies the hemisphere in which brain activities occur. CELM performs significantly better than all baselines on this task. Notably, the unimodal baseline with EEG features and text input captures frequency information but fails to localize brain activity spatially. In contrast, CELM and other baselines benefit from the learned channel embeddings of EEG foundation models, which enable better lateralization. ❐ Spectral dynamics. We study whether the generated report accurately captures frequency band behavior, which is a critical aspect of EEG interpretation. Interestingly, the unimodal baseline with handcrafted EEG features outperforms the multimodal baseline using linearly projected CBraMod embeddings, indicating that task-specific features can partially enable LLMs to infer specific EEG-related tasks. However, crafting such features is labor-intensive, motivating the use of foundation model embeddings. While naive linear projection fails to leverage these embeddings effectively, CELM achieves strong performance, demonstrating the importance of proper alignment strategies. ❐ Epileptiform abnormalities. We evaluate whether the generated report correctly identifies epileptiform activity. While all models perform reasonably well on this dimension, we note that epileptiform events are rare and highly imbalanced in the dataset.
5.4 Human Expert Evaluation Study
Figure 5: (a) Human expert evaluation study results, (b) Human expert preferential ranking for each report. (c) LLMs-as-a-judge results
To further strengthen our evaluation, we conducted an expert study in which six domain experts (clinicians and BME researchers) assessed report quality through an anonymized survey. Each expert evaluated 20 randomly sampled cases, with each case presenting the ground-truth clinical report alongside three anonymized generated reports from two different baseline categories and CELM. Experts rated each report on a 5-point scale (1: very poor to 5: very good) across five clinically relevant dimensions: (1) Clinical Correctness - accuracy compared to the ground truth, (2) Findings Coverage - coverage of key findings identified relative to ground truth, (3) Finding Accuracy — accuracy of attributes such as lateralization, frequency, and morphology, (4) Real-World Readiness — suitability for clinical use, and (5) Completeness — completeness of the report, whether all relevant aspects are covered. Results are summarized in Figure 5a, where CELM significantly outperforms both baselines across all five dimensions (Wilcoxon signed-rank test, p < 0.001 for all comparisons). Experts also provided their preferred rankings across the reports, as shown in Figure 5b, where CELM obtains a weighted-average rank of 1.29, with 98 votes for rank 1.
LLM-as-judge. Human expert evaluation is labor-intensive and difficult to scale. We conducted an LLM-as-judge evaluation of generated reports using the Qwen2.5-72B-Instruct-AWQ model (Figure 5c). Although closed LLMs (e.g., GPT-4, Claude) correlate well with human judgments, data-usage agreement constraints their use. We define three evaluation metrics, as in the expert study, and CELM consistently outperforms.
@@ 5.5 Alignment Module Ablation
Figure 6: (a) Report generation performance of different alignment module variants. (b) Training dynamics of each variant, including training loss, validation loss, and perplexity curves.
The alignment module is critical for ELMs, and naive linear projection has been shown to be insufficient. We ablate four alignment strategies: (1) Linear projector, projection via a single linear layer, (2) Perceiver projector, cross-attention with learnable queries followed by feedforward, and our proposed (3) SCA and (4) SCC projectors (Section 4.3). As shown in Figure 6, SCA achieves the best performance. Training dynamics (Figure 6b) reveal that Linear and Perceiver projectors overfit, with increasing validation loss despite decreasing training loss, whereas SCA and SCC exhibit stable convergence. This confirms confirming that modeling inter-epoch temporal dependencies prior to projection is essential for effective alignment.
@@ B.2 Evaluation Metrics
Table 4: Summary of LLM Baselines
LLM Family Context Length Gemma-3-1b-it Gemma 3 32K Gemma-3n-E2B-it Gemma 3 32K Gemma-3n-E4B-it Gemma 3 32K Gemma-3-4b-it Gemma 3 128K Medgemma-4b-it Gemma 128K Llama-3.2-1B-Instruct Llama-3.2 128K Llama-3.2-3B-Instruct Llama-3.2 128K Llama-3.1-8B-Instruct Llama-3.1 128K Meta-Llama-3-8B-Instruct Llama-3 8K Qwen3-1.7B Qwen 3 32K Qwen3-4B-Instruct-2507 Qwen 3 256K
@@ B.3 Baselines and Prompts
In this section, we provide additional details on the baseline models and the prompts used in our experiments. Table 4 summarizes all LLMs considered in our study along with their respective context lengths [^1] Context length information is extracted from their respective technical reports and Hugging Face https://huggingface.co/, further underscoring the practical challenges of directly conditioning LLMs on long-duration EEG recordings. Figures 8 and 9 present the prompts used for the unimodal text-only and unimodal text + EEG feature baselines, respectively. For clarity, we provide examples for certain prompt sections using bounding boxes.
@@
Patient history EEG description EEG SECTION DESCRIPTIONS [STANDARDIZED_SECTION_DESCRIPTIONS] e.g. EEG DESCRIPTION/DETAILS: Detailed narrative of EEG findings including background activity, sleep stages, physiologic variants, and abnormalities observed during the recording period. GUIDELINES Generate only the sections listed in **SECTIONS TO BE GENERATED**. Do NOT generate any additional sections. Do NOT repeat the same section more than once. Only generate the output in the JSON format and do not include any other text or explanation. OUTPUT FORMAT (STRICT)
@@
json {"report_sections": [ {"section_name": "Name of the section as given in SECTIONS TO BE GENERATED", "section_text": "Generated text for the section in string"}, ... ]} SECTIONS TO BE GENERATED
@@
Unimodal + Text + EEG Features Prompt EEG-DERIVED STATISTICS {"eeg_session_0":{"delta (0.5-4Hz) band power (dB)": [[-22.90, -21.60, ...]], "theta (4-8Hz) band power (dB)": [[-25.30, -24.80, ...]], "alpha (8-12Hz) band power (dB)": [[-27.50, -26.60, ...]], "beta (12-30Hz) band power (dB)": [[-30.60, -30.20, ...]], "gamma (30-80Hz) band power (dB)": [[-36.30, -36.10, ...]], EEG CHANNELS [’C3’, ’C4’, ’O1’, ’O2’, ’Cz’, ’F3’, ’F4’, ’F7’, ’F8’, ’Fz’, ’Fp1’, ’Fp2’, ’Fpz’, ’P3’, ’P4’, ’Pz’, ’T3’, ’T4’, ’T5’, ’T6’, ’A1’, ’A2’] You are an expert clinical neurophysiologist specializing in EEG interpretation and clinical report generation. TASK
@@
Patient history EEG description EEG Channels EEG-derived statistics (provided above) EEG SECTION DESCRIPTIONS [STANDARDIZED_SECTION_DESCRIPTIONS] e.g. EEG DESCRIPTION/DETAILS: Detailed narrative of EEG findings including background activity, sleep stages, physiologic variants, and abnormalities observed during the recording period. GUIDELINES Generate only the sections listed in **SECTIONS TO BE GENERATED**. Do NOT generate any additional sections. Do NOT repeat the same section more than once. Only generate the output in the JSON format and do not include any other text or explanation. OUTPUT FORMAT (STRICT)
@@
json {"report_sections": [ {"section_name": "Name of the section as given in SECTIONS TO BE GENERATED", "section_text": "Generated text for the section in string"}, ...]} SECTIONS TO BE GENERATED